OpenStudy (anonymous):

what is the area ,measured in square cenimeters of a triangle base is 9 height is 4 what is half of 9 ?

OpenStudy (anonymous):

\[\frac{9}{2}\]

OpenStudy (amistre64):

we can do area with determinates too ;)

OpenStudy (anonymous):

really?

OpenStudy (amistre64):

the determinate of a matrix is the area of the "rectangle"

OpenStudy (amistre64):

and a triangle is half that

OpenStudy (anonymous):

btw you do not need half of nine. what you need is \[A=\frac{base\times height}{2}=\frac{4\times 9}{2}=2\times 9\]]

OpenStudy (amistre64):



OpenStudy (amistre64):

det | a b | = area of "rectangle"

OpenStudy (anonymous):

hmm so what is a and b in this case?

OpenStudy (amistre64):

vectors

OpenStudy (anonymous):

ok . but what vectors are they in this example?

OpenStudy (amistre64):

b = 9,0 a = 0,4 if i read it right

OpenStudy (anonymous):

ok then i guess it must be right, since answer is 18

OpenStudy (amistre64):

\[{{9\ 0}\choose{0\ 4}}^{0}_{36}=36-0 = 36\]

OpenStudy (anonymous):

hell of a way to make a matrix!

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

@blanca half of 9 is 4.5, but your area is 18

OpenStudy (amistre64):

think of it as half the area of a rectangle the area of a rectangle is base x height so divide that in half